AnimeGhoul8863:

A quadratic equation is shown below: 25x2 + 10x + 1 = 0 Part A: Describe the solution(s) to the equation by just determining the radicand. Show your work. (5 points) Part B: Solve 4x2 − 4x + 1 = 0 by using an appropriate method. Show the steps of your work, and explain why you chose the method used. (5 points)

AnimeGhoul8863:

@Hero

Hero:

Part A just means to use \(b^2 - 4ac\) portion of the quadratic equation to determine if the quadratic is factorable.

AnimeGhoul8863:

can you explain what you mean

AnimeGhoul8863:

im confused

Hero:

The general form of a quadratic expression is \(ax^2 + bx + c\) The \(a,b,c\) values from that can be plugged in to the quadratic formula \(x = \dfrac{-b \pm \sqrt{b^2 - 4ac}}{2a}\)

AnimeGhoul8863:

ok

AnimeGhoul8863:

after you do dat den what e.e

Hero:

Do you think you can identify the \(a\) \(b\) and \(c\) values from the first quadratic?

AnimeGhoul8863:

ummmmmm ill try

AnimeGhoul8863:

a=25x^2 b=10x c=1

Hero:

\(a, b, c\) are only the coefficients. just the numbers not the variables

Hero:

a = 25, b = 10, c = 1

AnimeGhoul8863:

ok

AnimeGhoul8863:

Then what do we do once we have the numbers like that

Hero:

We put them in the \(b^2 - 4ac\) portion of the quadratic equation and see if we get a positive or negative value.

AnimeGhoul8863:

so 10^2 - 4(25)(1)?

Hero:

If \(b^2 - 4ac \ge 0\) then \(ax^2 + bx + c\) is factorable. If \(b^2 - 4ac < 0\) then \(ax^2 + bx + c\) is not factorable.

AnimeGhoul8863:

^^ok so do i do it this way Look above ur answer

AnimeGhoul8863:

@Hero ^^

Hero:

Yes, what was your result?

AnimeGhoul8863:

This always confused me sooo just a heads up xD

Hero:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Hero Yes, what was your result? \(\color{#0cbb34}{\text{End of Quote}}\)

AnimeGhoul8863:

idk i was waiting to see if it was yes or no so give me a sec plz

AnimeGhoul8863:

i got no clue i messed up some were or something cause im confused now

AnimeGhoul8863:

i got 0 as the answer

Hero:

Very good.

Hero:

So is Part A factorable?

AnimeGhoul8863:

No?

Hero:

Why not?

AnimeGhoul8863:

Bc its not > greater than zero its < zero

Hero:

What does it mean for \(b^2 - 4ac\) to be \(\ge\) to zero.

AnimeGhoul8863:

it can be factored

Hero:

Yes, correct.

AnimeGhoul8863:

\(\color{#0cbb34}{\text{Originally Posted by}}\) hero If \(b^2 - 4ac \ge 0\) then \(ax^2 + bx + c\) is factorable. If \(b^2 - 4ac < 0\) then \(ax^2 + bx + c\) is not factorable. \(\color{#0cbb34}{\text{End of Quote}}\) like this but i thought since its zero it wouldn't be factorable

AnimeGhoul8863:

@Hero so it is factorable

bm717:

@AnimeGhoul8863 Yes.

AnimeGhoul8863:

then what @Hero

AnimeGhoul8863:

@Hero You still there

AnimeGhoul8863:

@Shadow Can u pick up were hero left off

Hero:

So @AnimeGhoul8863, if it is factorable, then the next step would be to make an attempt to factor the expression.

AnimeGhoul8863:

What do you mean by that

AnimeGhoul8863:

@Hero

AnimeGhoul8863:

im confused again

Hero:

I've seen you factor quadratic expressions before. I'm surprised.

AnimeGhoul8863:

i think its bc i factor it when i see everything together but when its split up in multiple messages it confuses meh

AnimeGhoul8863:

@Hero you there you keep adding something and then leaves for like 10 mins

Hero:

Sorry about that

Hero:

We determined that \(25x^2 + 10x + 1\) was factorable and next step is to make an attempt to factor the expression. I've seen you factor in the past. Are you saying you don't have an approach to factoring this one?

AnimeGhoul8863:

25x^2+10x+1= (25x^2 +5x) + (5x+1) 5x(5x+1)+(5x+1) (5x+1)(5x+1) (5x+1)^2

AnimeGhoul8863:

^This @Hero

Hero:

Correct. The unique thing about when we get zero as the result of \(b^2 - 4ac\). When you factor \(ax^2 + bx + c\) you get a perfect square.

AnimeGhoul8863:

x'D @Hero Im sorry but can you help me word that we did the work but im confused how to put it all together

AnimeGhoul8863:

cause everything im trying to do to word this out isnt working

AnimeGhoul8863:

Can you put everything together for me so i can understand it and write it down please

AnimeGhoul8863:

@Hero

AnimeGhoul8863:

@Hero

AnimeGhoul8863:

All i need you to do is put everything together so i can write it down den were doneeeee

AnimeGhoul8863:

@Hero

AnimeGhoul8863:

e.e

Hero:

To summarize: The radicand is the number under the square root portion of the quadratic formula \(x = \dfrac{-b \pm \sqrt{b^2 - 4ac}}{2a}\) which can be determined by using the expression \(b^2 - 4ac\). If the value of \(b^2 - 4ac\) is \(> 0\), then \(ax^2 + bx + c\) is factorable. If the value of \(b^2 - 4ac\) is \(= 0\), then \(ax^2 + bx + c\) is a perfect square. If the value of \(b^2 - 4ac\) is \(< 0\), then \(ax^2 + bx + c\) is not factorable. In this case, we take the \(a, b, \text{ and } c\) values from \(25x^2 + 10x + 1\) and plug them in to the radicand. So in this case \(a = 25, b = 10, c = 1\). Plugging that into the radicand we get \(b^2 - 4ac = 10^2 - 4(25)(1) = 100 - 100 = 0\). So we know \(25x^2 + 10x + 1\) is a perfect square which we factor to get \((5x + 1)^2\). But we're not done because we still have to find the solution to the quadratic equation since the original expression was set equal to zero. So in fact, we have to find the solution to \((5x + 1)^2 = 0\). Once we have found the values for x, then will be finished.

AnimeGhoul8863:

im confused Whats The answer for Part A lets start there

AnimeGhoul8863:

we have done so much this is getting jumbled up

AnimeGhoul8863:

Part A

AnimeGhoul8863:

take the a,b, and c values from 25x2+10x+1 and plug them in to the radicand. a=25,b=10,c=1. Plugging that into the radicand we get b2−4ac=102−4(25)(1)=100−100=0. 25x2+10x+1 is a perfect square which we factor to get (5x+1)2.

Hero:

What I said above is most of Part A.

Hero:

All that was left was solving \((5x + 1)^2 = 0\) for \(x\).

AnimeGhoul8863:

ok stay here it will only take a sec

Hero:

And then afterwards you discover something about the solution to an equation with perfect square binomial/trinomial.

AnimeGhoul8863:

x=-1over 5

Hero:

Which is that there is only one solution. So the who point of part A was to get you to realize that the solution to \(25x^2 + 10x + 1 = 0\) has only one solution.

AnimeGhoul8863:

OMG its 10:00 PM AND WE HAVE TO DO PART B OMG T~T CAN WE TRY TO GO FAST WITH THIS PLZ

Hero:

So, what you can say here for Part A is that since the value of the radicand for \(25x^2 + 10x + 1\), is \( 0\), then there will only be one solution for the equation \(25x^2 + 10x + 1 = 0\)

Hero:

And then you say what that solution is.

Hero:

Sorry that took so long to type. I don't know why

AnimeGhoul8863:

Part A: : take the a,b, and c values from 25x2+10x+1 and plug them in to the radicand. a=25,b=10,c=1. Plugging that into the radicand we get b2−4ac=102−4(25)(1)=100−100=0. 25x2+10x+1 is a perfect square which we factor to get (5x+1)^2. Once we have (5x+1)^2 we solve for x x= -1 over 5 the solution to 25x2+10x+1=0 has only one solution ^^what i put

AnimeGhoul8863:

its ok lets just do part B

Hero:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Hero So, what you can say here for Part A is that since the value of the radicand for \(25x^2 + 10x + 1\), is \( 0\), then there will only be one solution for the equation \(25x^2 + 10x + 1 = 0\) \(\color{#0cbb34}{\text{End of Quote}}\) Put this but include the solution to the equation with it. And show your solution steps.

Hero:

You should show the steps you took to find the radicand and then show the steps you found to get the solution.

Hero:

Actually, for Part B, you use everything you learned in Part A to do that on your own.

AnimeGhoul8863:

Part A: : 25x2+10x+1, is 0, 25x2+10x+1=0 take the a,b, and c values from 25x2+10x+1 and plug them in to the radicand. a=25,b=10,c=1. Plugging that into the radicand we get <step b2−4ac=102−4(25)(1)=100−100=0. 25x2+10x+1 is a perfect square which we factor to get (5x+1)^2.<step Once we have (5x+1)^2 we solve for x <step x= -1 over 5 the solution to 25x2+10x+1=0 has only one solution ^^^the steps are there

AnimeGhoul8863:

OMG CAN YOU PLEASE JUST GIVE IT TOO ME IM DIEING HERE I CANT DO IT ON MY OWN THATS WHY IM HERE I JSUT WANNA GET THIS DONE SO PLEASE HELP ME GET THIS ALL DONE PLEASEEEEEEE

Hero:

Yeah, but you have to include what you understand about getting zero as the value of the radicand.

AnimeGhoul8863:

its almost 10:30 PM can you just bend the rules alittle bit u showed me how to do it just put it all together so i can put it down and be done PLEASEEEE

Hero:

@Shadow

AnimeGhoul8863:

@Hero Were did you gooooo

Hero:

The logic used to solve Part A is the same logic you use to solve Part B.

AnimeGhoul8863:

and what is that logic called again

AnimeGhoul8863:

quadratic method?

Hero:

Following the sequence of steps outlined here: 1. Determine the value of the radicand: The radicand is the number under the square root portion of the quadratic formula \(x = \dfrac{-b \pm \sqrt{b^2 - 4ac}}{2a}\) which can be determined by using the expression \(b^2 - 4ac\). 2. Determine from the value of the radicand whether the given quadratic expression is factorable: If the value of \(b^2 - 4ac\) is \(> 0\), then \(ax^2 + bx + c\) is factorable. If the value of \(b^2 - 4ac\) is \(= 0\), then \(ax^2 + bx + c\) is a perfect square. If the value of \(b^2 - 4ac\) is \(< 0\), then \(ax^2 + bx + c\) is not factorable. In this case, we take the \(a, b, \text{ and } c\) values from \(25x^2 + 10x + 1\) and plug them in to the radicand. So in this case \(a = 25, b = 10, c = 1\). Plugging that into the radicand we get \(b^2 - 4ac = 10^2 - 4(25)(1) = 100 - 100 = 0\). 3. Factor the expression: So we know \(25x^2 + 10x + 1\) is a perfect square which we factor to get \((5x + 1)^2\). 4. Find the solution to the original equation: But we're not done because we still have to find the solution to the quadratic equation since the original expression was set equal to zero. So in fact, we have to find the solution to \((5x + 1)^2 = 0\). Once we have found the values for x, then will be finished. \(\color{#0cbb34}{\text{End of Quote}}\)

AnimeGhoul8863:

Part A: the solution to 25x2+10x+1=0 has only one solution take the a,b, and c values from 25x2+10x+1 and plug them in to the radicand. a=25,b=10,c=1. Plugging that into the radicand we get b2−4ac=102−4(25)(1)=100−100=0. 25x2+10x+1 is a perfect square which we factor to get (5x+1)^2. Once we have (5x+1)^2 we solve for x x= -1 over 5 Part B: The method i used was quadratic method because it was easy to find the answer for part A with that method so i used the same method for part B. 4x2-4x+1=0 (22x2 - 4x) + 1 = 0 Factor 4x2-4x+1 4x2 - 2x - 2x - 1 2x * (2x-1) (2x-1) * (2x-1) Multiply (2x-1) by (2x-1) (2x - 1)2 = 0 (2x-1)2 = 0 solve for x x=1 over 2

AnimeGhoul8863:

^this is the best i can get it

Hero:

Okay, as long as you can show your steps. Some of your steps look a bit puzzling though.

AnimeGhoul8863:

Part A: (52x2 + 10x) + 1 = 0 Factor 25x2+10x+1 25 * 1 = 25 5 + 5 = 10 25x2 + 5x + 5x + 1 5x * (5x+1) (5x+1) * (5x+1) (5x + 1)2 = 0 x = -1/5

AnimeGhoul8863:

^better

Hero:

It would be better if you learned how to use \(LaTeX\)

AnimeGhoul8863:

Part A:the solution to 25x2+10x+1=0 has only one solution (52x2 + 10x) + 1 = 0 Factor 25x2+10x+1 25 * 1 = 25 5 + 5 = 10 25x2 + 5x + 5x + 1 5x * (5x+1) (5x+1) * (5x+1) (5x + 1)2 = 0 x = -1/5 Part B: The method i used was quadratic method because it was easy to find the answer for part A with that method so i used the same method for part B. 4x2-4x+1=0 (22x2 - 4x) + 1 = 0 Factor 4x2-4x+1 4x2 - 2x - 2x - 1 2x * (2x-1) (2x-1) * (2x-1) Multiply (2x-1) by (2x-1) (2x - 1)2 = 0 (2x-1)2 = 0 solve for x x=1 over 2

AnimeGhoul8863:

FOR THE LAST TIME I CANT

Hero:

You avoid using carets for exponents and other minor quirks

AnimeGhoul8863:

WHAT

MARC:

It looks like @AnimeGhoul8863 is using factorization method

AnimeGhoul8863:

Part A:the solution to 25x2+10x+1=0 has only one solution (52x2 + 10x) + 1 = 0 Factor 25x2+10x+1 25 * 1 = 25 5 + 5 = 10 25x2 + 5x + 5x + 1 5x * (5x+1) (5x+1) * (5x+1) (5x + 1)2 = 0 x = -1/5 Part B: The method i used was factorization method because it was easy to find the answer for part A with that method so i used the same method for part B. 4x2-4x+1=0 (22x2 - 4x) + 1 = 0 Factor 4x2-4x+1 4x2 - 2x - 2x - 1 2x * (2x-1) (2x-1) * (2x-1) Multiply (2x-1) by (2x-1) (2x - 1)2 = 0 (2x-1)2 = 0 solve for x x=1 over 2

Hero:

@MARC I did pick up on that much. Just some of the steps are not clear but also, she neglects use of carets for exponents.

AnimeGhoul8863:

HERO IF U WANNA ADD THEM ID BE HAPPPY FOR U TO DO THAT BUT IF ITS NOT A HUGE DEAL I SHOULD BE FINE

AnimeGhoul8863:

ITS 10:30 AT NIGHT I JUST WANNA FINISH THIS

MARC:

agree with u @Hero

Hero:

@AnimeGhoul8863 I think you finished it to the best of your ability. The solutions are not wrong. All you need to work on is your ability to explain what you're doing at each step. Hopefully, what I posted above can help.

AnimeGhoul8863:

you told me i didnt have to OMG im just gonna submit it the way it is i cant add anymore

AnimeGhoul8863:

Thank you hero for your help but this is what im submitting

AnimeGhoul8863:

Part A:the solution to 25x2+10x+1=0 has only one solution (52x2 + 10x) + 1 = 0 Factor 25x2+10x+1 25 * 1 = 25 5 + 5 = 10 25x2 + 5x + 5x + 1 5x * (5x+1) (5x+1) * (5x+1) (5x + 1)2 = 0 x = -1/5 Part B: The method i used was factorization method because it was easy to find the answer for part A with that method so i used the same method for part B. 4x2-4x+1=0 (22x2 - 4x) + 1 = 0 Factor 4x2-4x+1 4x2 - 2x - 2x - 1 2x * (2x-1) (2x-1) * (2x-1) Multiply (2x-1) by (2x-1) (2x - 1)2 = 0 (2x-1)2 = 0 solve for x x=1 over 2

AnimeGhoul8863:

^^what im submitting @Hero

AnimeGhoul8863:

Thank you for your help